OpenStudy (anonymous):

Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (mathstudent55):

150 people went Let x = the number of adults Then the number of students was 150 - x Adults paid $45 each, so the total amount paid by the adults was 45x Students paid $28.5 each, so the total amount paid by students was (150 - x)*28.5 The total paid is the amont paid by the adults plus the amount paid by the students, but we know this was $4770, so we can write: 45x + (150 - x)*28.5 = 4770 45x + 28.5(150 - x) = 4770 Solve for x to find the number of adults. Then subtract x from 150 to find the number of students.

OpenStudy (anonymous):

Let a and s be the number of adults and students respectively. Solve the following equations for a and s. Money is expressed as pennies. { a + s = 150, 4500 a + 2850 s = 477000 } { a = 30, s = 120 }